OpenStudy (anonymous):

the perimeter of a rectangle is 18ft. The length is 1 ft longer that the width. Let x be the length and y be the width. Stuck?

OpenStudy (phi):

length is 1 ft longer that the width x is length y is width replace "length" with x and "width" with y x is 1 ft longer than y is --> means = x = 1 ft longer than y 1 longer than ---> means 1+ x = 1 + y now the perimeter is the sum of the legs P = x+x+y + y = 2x + 2y x =(1+y) so we can replace x to get P = 2(1+y)+2y P is perimeter of a rectangle is 18ft. P = 18 so we now have 2(1+y)+2y=18 the next step is to solve for y (and then x using x= 1+y) distribute the 2 on the left side. can you do that ?

OpenStudy (anonymous):

Yes! Thank you for breaking it down for me!